Adams Furniture receives a special order for 10 sofas for a special price of $4,400. The direct materials and direct labor for each sofa are $170. In addition, supervision and other fixed overhead costs average $220 per sofa.

a1. What is the impact on operating income from accepting the special order?

a2. Based solely on a short-term financial analysis, should Adams accept the special order?

b1. If Adams is currently operating at full capacity, what would be the opportunity cost per unit for lost sales to regular customers if the special sales order is accepted and the selling price per unit on regular sales equals $500?

b2. Based solely on a short-term financial analysis, should Adams accept the special order if it is currently operating at full capacity?
Answer #1

Since there is excess capacity, impact on operating income = Revenue from special order - Variable cost of special order

= 4400 - 170*10

= $2,700

i.e. Operating income will increase by $2,700

a-2 Yes, should accept

b-1 Opportunity cost is the contribution margin lost on regular sales

= (500-170)*10

= $3,300

b-2 Effect on operating Income = 2700-3300

= -$600

No, should not accept